OpenStudy (anonymous):

y varies directly with x. When y = 12, x = 16. Find x when y = 24. 18 32 64

OpenStudy (callisto):

\[\frac{x_1}{y_1} = \frac{x_2}{y_2}\]\[\frac{16}{12} = \frac{x}{24}\]Can you solve it?

OpenStudy (anonymous):

i think so... would it be 32?

OpenStudy (callisto):

I think so :)

OpenStudy (anonymous):

thank you

OpenStudy (callisto):

Welcome~